22. A scientist runs a kinetic analysis and generates the following data 0 Reaction Progress She then proposes the following mechanism. Which rate law is consistent with these data and the proposed mechanism? Cl2 (g) + O2 (g) → Cl:O2 (g) CI202 (g) + N2 (g) → 2 ClNO (g) A. Rate Kobs Cl202]2 B. Rate = kobs[C12][02] C. Rate Kobs[Cl2]IO2][N2] D. Rate Kob2]N2] E. Rate Kobs CI O2]

Ans: C

Please explain why the answer is C. Thank you.
Answer #1

activation energy is larger for 2nd step
So, 2nd step is rate determining step

From 2nd step:
rate law is:
rate = k2[Cl2O2][N2]

since [Cl2O2] is intermediate, we will use step 1
using steady state approximation for step 1
rate forward = rate backward
k1[Cl2][O2] = k-1 [Cl2O2]
So,
[Cl2O2] = (k1/k-1)[Cl2][O2]

put this in rate law expression:
rate = k2[Cl2O2][N2]
rate = k2(k1/k-1)[Cl2][O2][N2]
let k2(k1/k-1) = kobs
so,
rate law is:
rate = kobs[Cl2][O2][N2]

Answer: C